OpenStudy (anonymous):

s^6/5 expression as radical

OpenStudy (mertsj):

Just remember: The top number is the power and the bottom number is the index of the radical.

OpenStudy (anonymous):

5/s^6 ???

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[\sqrt[5]{s^6}\]

OpenStudy (anonymous):

I have that but it is giving partial credit.

OpenStudy (mertsj):

Perhaps it wants you to simplify. Did it say that?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

\[\sqrt[5]{s^6}=\sqrt[5]{s^5\times s}=s \sqrt[5]{s}\]

OpenStudy (mertsj):

You might try that.

OpenStudy (anonymous):

Thanks

OpenStudy (mertsj):

yw. But it should have said in simplest form if it wanted you to simplify

OpenStudy (anonymous):

I agree.

OpenStudy (mertsj):

Did it like that answer?

OpenStudy (anonymous):

no, its has an extra box..... 5/s^6____ or S/5^s____. This is weird.

OpenStudy (mertsj):

Maybe I misunderstood the original problem. Was it this: \[\frac{s^6}{5}\]

OpenStudy (mertsj):

Or this: \[s ^{\frac{6}{5}}\]

OpenStudy (anonymous):

no, you had it right the first time. \[S ^\frac{ 6 }{5}\]

OpenStudy (mertsj):

So it's that second one.

OpenStudy (anonymous):

yeah, sorry late post